OpenStudy (anonymous):

simplify; the square root of 50x^(2)

OpenStudy (anonymous):

\[\sqrt{50x^{2}} = \sqrt{25} * \sqrt{x ^{2}} * \sqrt{2} = 5*x*\sqrt{2} = 5x \sqrt{2}\]